A random sample of 11 observations was taken from normal population. The sample mean and

standard deviation are 74.5 and 9 accordingly. Can we infer at 5% significance level that the

population mean is greater than 70?

5. Repeat number 4 with assuming the population standard deviation = 9


1
Expert's answer
2022-05-11T12:09:32-0400

4. The following null and alternative hypotheses need to be tested:

"H_0:\mu\le70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=10" degrees of freedom, and the critical value for a right-tailed test is "t_c = 1.812461."

The rejection region for this right-tailed test is "R = \{t: t> 1.812461\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{74.5-70}{9/\sqrt{11}}\approx1.6583"


Since it is observed that "t=1.6583< 1.812461=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=10" degrees of freedom, "t=1.6583" is "p=0.064124," and since "p=0.064124>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is greater than 70, at the "\alpha = 0.05" significance level.


5. The following null and alternative hypotheses need to be tested:

"H_0:\mu\le70"

"H_a:\mu>70"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c = 1.6449."

The rejection region for this right-tailed test is "R = \{z: z> 1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{74.5-70}{9/\sqrt{11}}\approx1.6583"


Since it is observed that "z=1.6583> 1.6449=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed is "p=P(Z>1.6583)=0.048628," and since "p=0.048628<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 70, at the "\alpha = 0.05" significance level.